Consider the shuttle carrying the astronauts is dropped from 100ft. The position 𝑠 of the shuttle is given by the following function as a function of time: 𝑠(𝑑) = βˆ’16𝑑 2 + 100

Consider the interval [0, c] where 0 is the time when the shuttle drops, and c is the time when the shuttle hits the surface of the Earth. Find the height or position of the shuttle when time is: 0, 0.5, 1, 1.5, 2, and 2.5 seconds. What is the domain of this function? Also, sketch the graph of the position s.Β 


1
Expert's answer
2021-09-27T12:25:33-0400
"s(t)=-16t^2+100"

"s(0)=-16(0)^2+100=100 (ft)"

"s(0.5)=-16(0.5)^2+100=96 (ft)"

"s(1)=-16(1)^2+100=84 (ft)"

"s(1.5)=-16(1.5)^2+100=64 (ft)"

"s(2)=-16(2)^2+100=36 (ft)"

"s(2.5)=-16(2.5)^2+100=0 (ft)"



"Domain:[0,2.5]"